this is a list of ingredients to make 16 brownies. use this list to answer both questions:
2/3 cup butter
5 squares (1 ounce ea) unsweetened chocolate
1 1/2 cups of sugar
2 teaspoons vanilla
2 eggs
1 cup flour
1. How much sugar is needed to make 8 brownies?
2. What is the greatest # of brownies that can be made if baker only has 1 cup of butter

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

8/16=1.5/X CROSS MULTIPLY
8X-1.5*16
8X=24
X=3 CUPS OF SUGAR ARE REQUIRED.
--------------------------------------
(2/3)/1=16/X
2X/3=16
2X=3*16
2X=48
X=48/2
X=24 COOKIES CAN BE MADE WITH I CUP OF BUTTER.

Answer by stanbon(75887)

1 1/2 cups of sugar for 16 cookies
--------------
1. How much sugar is needed to make 8 brownies?
(1/2)(3/2) = 3/4 cup of sugar for 8 cookies
-----------------------
2. What is the greatest # of brownies that can be made if baker only has 1 cup of butter
x/1 = 16/(3/2)
x = 32/3 = 10 1/3 cookies from 1 lb of butter.
================
Cheers,
Stan H.